Randomized Algorithms - Course Review

K. Subramani

LCSEE, West Virginia University,

Morgantown, WV {ksmani@csee.wvu.edu}

April 24, 2012

1. Are you comfortable with the pace of the class? (a) Yes.

(b) No.

(c) Most of the time.

2. Is the course challenging and interesting?

(a) Yes.

(b) No.

(c) Sometimes.

(d) Most of the time.

3. How much time do you spend on this subject outside of class?

(a) None.

(b) 1-2 hours per week.

(c) 4-5 hours per week.

4. Are you learning new material in this course? (a) A lot.

(b) Some.

(c) No.

5. Are your critical thinking skills improving through this course?

(a) Yes, significantly.

(b) Somewhat.

(c) No.

6. How is the workload vis-a-vis other courses? (a) About the same.

(b) Less than other courses.

(c) More than other courses.

7. Is the lecturer effective at communicating ideas? (a) Always.

(b) Most of the time.

(c) Sometimes.

(d) Never.

8. Is the lecturer well-organized?

(a) Always.

(b) Never.

(c) Most of the time.

(d) Sometimes.

9. Have you met the instructor during office hours? (a) Occasionally.

(b) Never.

10. Has the instructor been helpful to you, outside class?

(a) Yes.

(b) No.

11. How would you rate this instructor vis-a-vis other instructors that you have had?

(a) Better than other instructors.

(b) As good as other instructors.

(c) Worse than other instructors.

12. How would you rate this course vis-a-vis other courses that you have taken?

(a) Better than other courses.

(b) As good as other courses.

(c) Worse than other courses.

13. Should this course become a permanent offering in the department?

(a) Yes.

(b) No.

14. Does the instructor show mastery over the material?

(a) Yes.

(b) No.

15. Is the instructor enthusiastic about the material? (a) Always.

(b) Most of the time.

(c) Sometimes.

(d) Never.

16. Would you take another course from this instructor?

(a) Definitely.

(b) Highly likely.

(c) Unlikely.

(d) Only if hell freezes over.

17. What aspects of the course would you like to maintain?

18. What aspects of the course would you like to change and in what way?